### v1

# Hausman’s Consistency Test and an Internal Bias Diagnostic: A Short Note Regarding Linear Mixed Models

Abstract.
The Hausman specification test detects inconsistency in mixed model estimators of fixed effects by comparing the original model with an alternative specification in which random effects are treated as fixed. This note illustrates a bias diagnostic from the statistical literature as a complement to the Hausman test. The diagnostic can provide internal estimates of parameter-specific bias in the mixed model estimators without requiring the second, fixed-effects-only model to be fit. We apply the diagnostic to a panel data analysis as well as to a Value-Added Model (VAM) for teacher evaluation.
Key words and phrases:
linear mixed model, misspecification, permutation test, stochastic model matrix
1. Introduction
Consider a linear mixed model:
(1) ${\mathbf{y}}={\mathbf{X}}\boldsymbol{\beta}+{\mathbf{Z}}\boldsymbol{\eta}+\boldsymbol{\epsilon},$
with $\boldsymbol{\eta}\sim N(\boldsymbol{0},{\mathbf{G}})$ independent of $\boldsymbol{\epsilon}\sim N(\boldsymbol{0},{\mathbf{R}})$. For the resulting estimator of $\boldsymbol{\beta}$ to be consistent, it is necessary that the random effects $\boldsymbol{\eta}$ are uncorrelated with the regressors in ${\mathbf{X}}$ (Hausman & Taylor, 1981a). The Hausman test (Hausman, 1978) provides a classical tool to assess the consistency of this mixed model estimator by comparing the random-$\boldsymbol{\eta}$ effects (RE) and fixed-$\boldsymbol{\eta}$ effects (FE) estimators, the latter requiring a re-fit of the model that treats $\boldsymbol{\eta}$ as an additional vector of fixed effects. Under the null hypothesis, both RE and FE estimators are consistent, resulting in only minor differences between them. Under the alternative, only the FE estimator is consistent. A large discrepancy suggests that the RE estimator is inconsistent, indicating a violation of the orthogonality assumptions or other modeling conditions (Hausman, 1978; Hausman & Taylor, 1981b).
The Hausman test considers consistency: it evaluates whether the RE estimator converges to the true parameter values as the sample size tends to infinity. By contrast, bias refers to the systematic difference between the expected value of an estimator and the true value, at any given sample size. While these are distinct concepts – it is possible for an estimator to be asymptotically consistent yet still exhibit finite-sample bias – similar sources of model misspecification (such as violations of exogeneity) often affect both consistency and bias. To this end, it is reasonable to use a bias diagnostic alongside the Hausman test when considering potential model misspecification.
We illustrate an application of the work of Karl & Zimmerman (2021), who derive an internal diagnostic for bias in linear mixed models when the random-effects design matrix ${\mathbf{Z}}$ may be correlated with the random effects $\boldsymbol{\eta}$. Their diagnostic produces a plug-in bias estimate, $\hat{\boldsymbol{\nu_{k}}}^{\prime}\hat{\boldsymbol{\eta}}$, and a permutation-based p-value, all without requiring a second, fixed-effects-only model fit. This feature is especially useful in complex applications (such as the value-added model in Section 4), where treating $\boldsymbol{\eta}$ as fixed would be impractical due to insufficient degrees of freedom or lack of off-the-shelf software availability. For further theoretical details on this bias formula and its permutation scheme, see Karl & Zimmerman (2021).
In this note, we demonstrate how the bias diagnostic can be deployed alongside the well-known Hausman specification test. While the Hausman test evaluates consistency by comparing RE and FE estimators, the internal bias diagnostic measures finite-sample bias due to correlations between ${\mathbf{Z}}$ and $\boldsymbol{\eta}$. Together, these tools provide a more comprehensive check of potential misspecification in linear mixed models, and we illustrate their complementary roles using both a panel-data example and a value-added model.
In Section 2, we present both the Hausman test and the bias diagnostic. Section 3 applies these methods to a gasoline consumption dataset using the R packages plm and mixedbiastest, and Section 4 illustrates the bias diagnostic in a value-added modeling (VAM) context using the R package GPvam.
2. Hausman and Bias Diagnostics
2.1. The Hausman Test
The Hausman test (Hausman, 1978) compares the FE and RE estimators for $\boldsymbol{\beta}$, $\hat{\boldsymbol{\beta}}_{\text{FE}}$ and $\hat{\boldsymbol{\beta}}_{\text{RE}}$. Under the null hypothesis, both estimators are consistent. Letting ${\mathbf{M}}^{-}$ represent a generalized inverse of ${\mathbf{M}}$, the test statistic
$H=(\hat{\boldsymbol{\beta}}_{RE}-\hat{\boldsymbol{\beta}}_{FE})^{\prime}[\operatorname{Var}(\hat{\boldsymbol{\beta}}_{FE})-\operatorname{Var}(\hat{\boldsymbol{\beta}}_{RE})]^{-}(\hat{\boldsymbol{\beta}}_{RE}-\hat{\boldsymbol{\beta}}_{FE})$
is $\chi^{2}(m)$ under $H_{0}$, where $m$ is the rank of $\operatorname{Var}(\hat{\boldsymbol{\beta}}_{FE})-\operatorname{Var}(\hat{\boldsymbol{\beta}}_{RE})$ (Hausman & Taylor, 1981a). A significant $H$ indicates that the RE assumptions do not hold, and the RE estimator is inconsistent.
2.2. Bias Diagnostic
The RE estimator is
$\hat{\boldsymbol{\beta}}_{\text{RE}}=({\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{X}})^{-}{\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{y}},$
with $\hat{{\mathbf{V}}}={\mathbf{Z}}\hat{{\mathbf{G}}}{\mathbf{Z}}^{\prime}+\hat{\mathbf{R}}$. Under the assumption that ${\mathbf{X}}$ is fixed and ${\mathbf{Z}}$ is stochastic, Karl & Zimmerman (2021) show
$\operatorname{Bias}(\mathbf{k}^{\prime}\hat{\boldsymbol{\beta}}_{\text{RE}})=\mathbf{E}[\hat{\boldsymbol{\nu_{k}}}^{\prime}\boldsymbol{\eta}]$
where
$\hat{\boldsymbol{\nu_{k}}}=\mathbf{k}^{\prime}({\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{X}})^{-}{\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{Z}}$
and that this bias is zero if any of the following three conditions are true: (1) ${\mathbf{X}}^{\prime}\mathbf{R}^{-1}{\mathbf{Z}}=\boldsymbol{0}$, (2) if $\boldsymbol{\eta}$ and ${\mathbf{Z}}$ are independent, or (3) if $\boldsymbol{\eta}$ is treated as fixed (producing the FE estimator). The third condition indicates that the FE estimator remains unbiased in the presence of model misspecification that biases the RE estimator. Thus the internal bias estimate often approximates the difference between the RE and FE estimators.
The plug-in quantity $\hat{\boldsymbol{\nu_{k}}}^{\prime}\hat{\boldsymbol{\eta}}$ serves as an internal bias estimate for the RE estimator. A vector permutation function $\boldsymbol{\pi}(\cdot)$ is applied repeatedly to calculate draws $\hat{\boldsymbol{\nu_{k}}}^{\prime}\boldsymbol{\pi}(\hat{\boldsymbol{\eta}})$ from a reference distribution to assess if the observed bias is unusually large under the null hypothesis that $\boldsymbol{\eta}$ is independent of all other model components. The bias p-value is calculated as the proportion of reference values that are larger in absolute value than that of the internal plug-in bias estimate. This test has been implemented in the R package mixedbiastest (Karl, 2024): the package accepts mixed model objects from the popular lme4 package (Bates et al., 2015) and produces bias diagnostics for each parameter of $\boldsymbol{\beta}$ as well as custom contrasts. mixedbiastest currently only accepts lme4 models with diagonal random effects covariance matrices $({\mathbf{G}})$, but this matches the structure of most panel data applications and could be generalized as described by Karl & Zimmerman (2021).
3. Gasoline Consumption Data Application
We illustrate both tests using the Gasoline dataset from the R package plm, which contains panel data on gasoline consumption (Baltagi & Griffin, 1983; Croissant & Millo, 2008). Listing 1 presents the R code to run the Hausman test and to load libraries for the bias test.
1# Load required libraries
2library(plm)
3library(lme4)
4library(mixedbiastest)
5# Load the Gasoline dataset
6data("Gasoline", package = "plm")
7# Define the model formula
8form <- lgaspcar ~ lincomep + lrpmg + lcarpcap
9# Within (FE) model
10wi <- plm(form, data = Gasoline, model = "within")
11# Random effects model
12re <- plm(form, data = Gasoline, model = "random")
13# Hausman test
14htest <- phtest(wi, re)
15# Results
16print(htest)
Listing 1:  Hausman Test for the Gasoline Dataset
Hausman Test
chisq = 302.8, df = 3, p-value < 2.2e-16
The significant result rejects the null hypothesis that the RE estimator is consistent.
Next, we fit the mixed model with lme4 and run the bias diagnostic (Listing 2). For comparison of parameter estimates, we also fit the corresponding fixed effects model, although this is not required by the bias test
1# mixed model
2random_model <- lmer(lgaspcar ~ lincomep + lrpmg + lcarpcap + (1 | country),
3 data = Gasoline)
4# run the bias test
5res<-mixedbiastest(random_model, n_permutations = 1e6, verbose = FALSE)
6# treat country indicators as fixed effects
7fixed_model <- lm(lgaspcar ~ lincomep + lrpmg + lcarpcap + country,
8 data = Gasoline)
9print(random_model)
10print(fixed_model)
11print(res)
Listing 2:  Bias Test for the Gasoline Dataset
Figure 1.  Graphical Summary of Bias Diagnostics for Gasoline Data. The histograms show the reference distributions, $\hat{\boldsymbol{\nu_{k}}}^{\prime}\boldsymbol{\pi}(\hat{\boldsymbol{\eta}}),$ and the dashed red lines show the internal bias estimates, $\hat{\boldsymbol{\nu_{k}}}^{\prime}\hat{\boldsymbol{\eta}}$.
Table 1.  Gasoline Data: FE, RE Estimates, RE - FE difference, Internal Bias Estimates, and Bias p-value (rounded)
Parameter FE Est RE Est (RE - FE) Bias Estimate Bias p-value
(Intercept) 2.29 2.15 -0.14 -0.17 0.1048
lincomep 0.66 0.59 -0.07 -0.04 0.1596
lrpmg -0.32 -0.37 -0.05 -0.04 0.0008
lcarpcap -0.64 -0.62 0.02 0.01 0.1975
The graphical summaries of the bias estimates and reference distributions appear in Figure 1. The results from Listing 2 are combined into Table 1. The Hausman test indicated inconsistency of the RE estimator, and bias diagnostic now suggests that lrpmg is significantly biased downward. Note the similarity of the internal bias estimates from the RE model with the RE-FE differences in Table 1.
4. Value-added Model (VAM) Results with GPvam
Linear mixed models are used for value-added assessment of teachers based on the standardized test scores of their students (Morganstein & Wasserstein, 2014). There are many variations of these models, but one particular instance is the complete persistence (CP) VAM (Ballou et al., 2004; Mariano et al., 2010; Karl et al., 2013b). In the CP model, intra-student correlation is modeled with off-diagonal entries in the error covariance matrix (${\mathbf{R}}$). Classroom-heterogeneity – the vector of individual “teacher effects” (Lockwood et al., 2007) – is modeled with random effects ($\boldsymbol{\eta}$). A different variance component is fit for teachers in different grades, producing a diagonal covariance matrix ${\mathbf{G}}$ with the number of unique elements equal to the number of grades in the study. The random effects model matrix ${\mathbf{Z}}$ tracks the assignment of students to classrooms, and also attributes student scores after the first year to their previous years’ teachers, producing a “multiple member” structure for ${\mathbf{Z}}$ (Browne et al., 2001). The nonrandom assignment of students to classrooms creates the potential for biasing (Lockwood & McCaffrey, 2007) according to the mechanism described by Karl & Zimmerman (2021). Calculating the Hausman test statistic to test consistency for this example would require customized programming to fit a FE model that treats the teacher effects as fixed while still utilizing the multiple membership structure of ${\mathbf{Z}}$ and the block diagonal structure of the error covariance matrix, $\mathbf{R}$. By contrast, the bias test only requires the estimates and model matrices from the fitted mixed model.
The R package GPvam (Karl et al., 2012, 2013b) in Listing 3 fits the CP model and has a function to calculate the bias diagnostic. Note that GPvam requires the provided R data object to contain columns named “teacher”, “year”, and “student” so that the ${\mathbf{Z}}$, ${\mathbf{G}}$, and $\mathbf{R}$ matrices can be constructed. This is implicit in the function call, and the formula argument to GPvam only constructs the fixed effects structure ${\mathbf{X}}$ and $\boldsymbol{\beta}$. We examine the bias results for the application of the CP model to math test scores from a large urban elementary school dataset with 2834 students from grades 4 to 6. This is the same data set used by Karl et al. (2013a).
1library(GPvam)
2result.cp<-GPvam(data, persistence = "CP",
3 formula(~as.factor(Race_Ethnicity)+as.factor(year)+as.factor(Gender)+0),
4 hessian = TRUE,tol1=1e-9)
5# Perform the bias test on all fixed effects
6bias.test.custom(result.cp)
7# Perform the bias test on a contrast: White - Hispanic
8bias.test.custom(result.cp,n_perms=1e6, k_vectors = list(c(0,0,0,-1,1,0,0,0)))
Listing 3:  Fitting a CP model with GPvam and performing the bias diagnostic
Table 2.  CP Model Fixed Effects Estimates, Standard Errors, Internal Bias Estimates, and Bias p-values
Fixed Effect Estimate Std. Error Bias Estimate Bias p-value
Black 24.2894 0.1612 0.0062 0.7430
Am Indian/Alaska 24.3683 0.3228 -0.0364 0.1570
Asian/Pac Island 25.7910 0.1863 0.0506 0.0105
Hispanic 24.3989 0.1058 -0.0691 0.0004
White 25.4836 0.0894 0.0597 0.0000
year=1 - - - -
year=2 0.9831 0.0600 0.0023 0.5440
year=3 1.9738 0.0876 0.0012 0.7530
Gender=F - - - -
Gender=M 0.0552 0.0819 0.0036 0.4990
The CP estimates and bias diagnostics for the fixed effects appear in Table 2. The Hispanic parameter estimate is significantly downward-biased, while the White and Asian/Pac Island parameter estimates are significantly upward-biased. The contrast White - Hispanic reveals even stronger bias with a point estimate of $0.1287$ and a p-value of 0 (none of the one million permutations exceeded the test statistic). Although the bias is statistically significant, its magnitude may not be practically significant for this particular application. Nevertheless, this example illustrates how the diagnostic can facilitate the investigation of potential bias due to nonrandom allocation of random effects on a case-by-case basis.
5. Conclusion
We have illustrated how the bias test of Karl & Zimmerman (2021) can be used to detect instances where the mixed model (1) produces biased estimators for fixed effects. Besides considering bias rather than consistency, this test is narrower than the Hausman test in the sense that it focuses on bias due to dependence between the random effects and their corresponding model matrix in situations where the fixed and random effects model matrices are not orthogonal (with respect to the inverse of the error covariance matrix). In contrast, the Hausman test is sensitive to any potential source of inconsistency in the random-$\boldsymbol{\eta}$ estimator of the fixed effects, provided the fixed-$\boldsymbol{\eta}$ estimator is consistent. Nevertheless, the bias diagnostic can rely solely on the fitted random-$\boldsymbol{\eta}$ model, generate summaries for each fixed effect or contrast, and use a permutation-based reference distribution. These features make it a potential complement to the Hausman test when assessing potential misspecification in linear mixed models.