### v3

# Parameter-Specific Bias Diagnostics in Random-Effects Panel Data Models

Abstract.
The Hausman specification test detects inconsistency of the random-effects estimator by comparing it with an alternative fixed-effects estimator. This note shows how a recently proposed bias diagnostic for linear mixed models can complement this test in random-effects panel-data applications. The diagnostic delivers parameter-specific internal estimates of finite-sample bias of the random-effects estimator, together with permutation-based $p$-values, from a single fitted random-effects model. We illustrate its use in a gasoline-demand panel and in a value-added model for teacher evaluation, using publicly available R packages, and we discuss how the resulting bias summaries can be incorporated into routine practice.
Key words and phrases:
linear mixed model, Hausman test, correlated random effects, permutation test
1. Introduction
Consider a linear mixed model:
(1) ${\mathbf{y}}={\mathbf{X}}\boldsymbol{\beta}+{\mathbf{Z}}\boldsymbol{\eta}+\boldsymbol{\epsilon},$
with $\boldsymbol{\eta}\sim N(\boldsymbol{0},{\mathbf{G}})$ independent of $\boldsymbol{\epsilon}\sim N(\boldsymbol{0},\mathbf{R})$. Here $\mathbf{y}$} is an $n\times 1$ response vector, $\mathbf{X}$} is an $n\times p$ fixed-effects design matrix with associated coefficient vector $\boldsymbol{\beta}$, and $\mathbf{Z}$} is an $n\times m$ random-effects design matrix with random-effects vector $\boldsymbol{\eta}$. The matrices $\mathbf{G}$} and $\mathbf{R}$ denote the $m\times m$ random-effects covariance matrix and the $n\times n$ error covariance matrix, respectively.
In random-effects panel-data applications, consistency of the RE estimator requires the unobserved individual effects to be uncorrelated (or, more generally, mean independent) with the regressors (the usual exogeneity condition; see, e.g., Hausman & Taylor, 1981a). The Hausman test (Hausman, 1978) provides a classical tool to assess the consistency of this mixed model estimator by comparing the random-effects (RE) and fixed-effects (FE) estimators. In its original formulation, this involves re-fitting the model with $\boldsymbol{\eta}$ treated as an additional vector of fixed effects, although correlated random-effects (CRE) implementations in the spirit of Mundlak (1978) and Wooldridge (2010) deliver Hausman-type specification tests based on a single augmented RE fit; see Section 2. Under the null hypothesis, both RE and FE estimators are consistent, resulting in only minor differences between them. Under the alternative, only the FE estimator is consistent. A large discrepancy suggests that the RE estimator is inconsistent, indicating a violation of the orthogonality assumptions or other modeling conditions (Hausman, 1978; Hausman & Taylor, 1981b).
The Hausman test considers consistency: it evaluates whether the RE estimator converges to the true parameter values as the sample size tends to infinity. By contrast, bias refers to the systematic difference between the expected value of an estimator and the true value, at any given sample size. While these are distinct concepts – it is possible for an estimator to be asymptotically consistent yet still exhibit finite-sample bias – similar sources of model misspecification (such as violations of exogeneity) often affect both consistency and bias. To this end, it is reasonable to use a bias diagnostic alongside the Hausman test when considering potential model misspecification. In what follows, we deliberately keep these roles separate: the Hausman test is used as a classical large-sample specification test for the RE estimator, whereas the diagnostic of Karl & Zimmerman (2021) is used as a finite-sample tool to quantify the magnitude and direction of parameter-specific bias in the same RE fit.
We illustrate an application of the work of Karl & Zimmerman (2021), who derive an internal diagnostic for bias in linear mixed models when, conditional on the fixed-effects design matrix $\mathbf{X}$}, the random-effects design matrix $\mathbf{Z}$} may be stochastically dependent on the random effects $\boldsymbol{\eta}$. Their diagnostic produces a plug-in bias estimate, $\hat{\boldsymbol{\nu_{k}}}^{\prime}\hat{\boldsymbol{\eta}}$, and a permutation-based $p$-value, all without requiring a second, fixed-effects-only model fit. This feature is especially useful in complex applications (such as the value-added model in Section 4), where treating $\boldsymbol{\eta}$ as fixed would be impractical due to insufficient degrees of freedom or lack of off-the-shelf software availability. For further theoretical details on this bias formula and its permutation scheme, see Karl & Zimmerman (2021).
In this note, we demonstrate how the bias diagnostic can be deployed alongside the well-known Hausman specification test. While the Hausman test evaluates consistency by comparing RE and FE estimators, the internal bias diagnostic measures finite-sample bias due to correlations between $\mathbf{Z}$} and $\boldsymbol{\eta}$. Together, these tools provide a more comprehensive check of potential misspecification in linear mixed models, and we illustrate their complementary roles using both a panel-data example and a value-added model. Our aim is not to propose a new alternative to the Hausman or Mundlak–Wooldridge tests, but rather to show how an existing linear mixed model bias diagnostic can supplement them by providing coefficient- or contrast-specific information obtained from a single fitted mixed model.
In Section 2, we present both the Hausman test and the bias diagnostic. Section 3 applies these methods to a gasoline consumption dataset using the R packages plm and mixedbiastest, and Section 4 illustrates the bias diagnostic in a value-added modeling (VAM) context using the R package GPvam.
2. Hausman and Bias Diagnostics
2.1. The Hausman Test
The Hausman test (Hausman, 1978) compares the FE and RE estimators for $\boldsymbol{\beta}$, $\hat{\boldsymbol{\beta}}_{\text{FE}}$ and $\hat{\boldsymbol{\beta}}_{\text{RE}}$. Under the null hypothesis, both estimators are consistent. Letting $\mathbf{M}$$^{-}$ represent a generalized inverse of a matrix $\mathbf{M}$}, the test statistic
$H=(\hat{\boldsymbol{\beta}}_{\text{RE}}-\hat{\boldsymbol{\beta}}_{\text{FE}})^{\prime}\bigl[\operatorname{Var}(\hat{\boldsymbol{\beta}}_{\text{FE}})-\operatorname{Var}(\hat{\boldsymbol{\beta}}_{\text{RE}})\bigr]^{-}(\hat{\boldsymbol{\beta}}_{\text{RE}}-\hat{\boldsymbol{\beta}}_{\text{FE}})$
is $\chi^{2}(m)$ under $H_{0}$, where $m$ is the rank of $\operatorname{Var}(\hat{\boldsymbol{\beta}}_{\text{FE}})-\operatorname{Var}(\hat{\boldsymbol{\beta}}_{\text{RE}})$ (Hausman & Taylor, 1981a). A significant $H$ indicates that the RE assumptions do not hold, and the RE estimator is inconsistent. In the canonical panel-data setting, this statistic is derived under the assumption of homoskedastic, serially uncorrelated errors, that is, $\mathbf{R}=\sigma^{2}\mathbf{I}_{n}$. In the linear mixed model formulation we write the model with a general error covariance matrix $\mathbf{R}$; the gasoline example below maintains the standard homoskedastic assumption, whereas the value-added model in Section 4 uses a block-diagonal $\mathbf{R}$.
2.2. Bias Diagnostic
Let $\hat{{\mathbf{G}}}$ and $\hat{\mathbf{R}}$ denote estimates of $\mathbf{G}$} and $\mathbf{R}$, and define $\hat{{\mathbf{V}}}={\mathbf{Z}}\hat{{\mathbf{G}}}{\mathbf{Z}}^{\prime}+\hat{\mathbf{R}}$. The empirical random-effects (RE) estimator of $\boldsymbol{\beta}$ is
(2) $\hat{\boldsymbol{\beta}}_{\text{RE}}=({\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{X}})^{-}{\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{y}}.$
Following Karl & Zimmerman (2021), we treat $\mathbf{X}$} as fixed and $\mathbf{Z}$} as stochastic. For any $p\times 1$ vector $\mathbf{k}$ such that $\mathbf{k}^{\prime}\boldsymbol{\beta}$ is an estimable linear combination of the fixed effects (for example, a single regression coefficient $\beta_{j}$ or a contrast such as $\beta_{j}-\beta_{\ell}$), define
(3) $\hat{\boldsymbol{\nu_{k}}}=\mathbf{k}^{\prime}({\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{X}})^{-}{\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{Z}}.$
Karl & Zimmerman (2021) show that, under regularity conditions,
(4) $\operatorname{Bias}(\mathbf{k}^{\prime}\hat{\boldsymbol{\beta}}_{\text{RE}})=\mathbf{E}[\hat{\boldsymbol{\nu_{k}}}^{\prime}\boldsymbol{\eta}].$
The bias of the random-effects estimator $\mathbf{k}^{\prime}\hat{\boldsymbol{\beta}}_{\text{RE}}$ is zero if either (i) ${\mathbf{X}}^{\prime}\mathbf{R}^{-1}{\mathbf{Z}}=\boldsymbol{0}$ or (ii) the random-effects vector $\boldsymbol{\eta}$ is independent of $\mathbf{Z}$}. A third way to avoid bias in practice is to treat $\boldsymbol{\eta}$ as fixed (producing the FE estimator), in which case the corresponding fixed-effects estimator of $\mathbf{k}^{\prime}\boldsymbol{\beta}$ remains unbiased.
[…]
2.3. Relation to correlated random effects and Mundlak–Wooldridge tests
In the panel-data literature, a common way to test the random-effects specification is through correlated random effects (CRE) or “Mundlak” type augmentations of the RE model. Suppose we have the panel-data setup
$y_{it}=x_{it}^{\prime}\beta+c_{i}+\varepsilon_{it},$
with individual effect $c_{i}$ and idiosyncratic error $\varepsilon_{it}$. Under the usual random-effects assumptions, including strict exogeneity and $E[c_{i}\mid x_{i1},\ldots,x_{iT_{i}}]=0$ (often summarized as $\operatorname{Cov}(x_{it},c_{i})=0$), the random-effects estimator is consistent. Mundlak (1978) models potential correlation between $x_{it}$ and $c_{i}$ by including the group means $\bar{x}_{i}$ in the regression. Wooldridge (2010) shows that a simple Hausman-type alternative is to estimate the augmented random-effects model
$y_{it}=x_{it}^{\prime}\beta+\bar{x}_{i}^{\prime}\gamma+c_{i}+\varepsilon_{it},$
and apply a (robust) Wald test of $H_{0}:\gamma=0$. This CRE test uses only an augmented RE fit but is asymptotically equivalent to the traditional FE–RE Hausman test when the idiosyncratic errors are homoskedastic and serially uncorrelated (Hausman & Taylor, 1981a). Robust implementations based on covariance estimators that are consistent under general forms of heteroskedasticity and serial correlation are widely used in applied work; see, e.g., Wooldridge (2010).
In terms of the linear mixed model (1), stacking the panel outcomes $\{y_{it}\}$ into $\mathbf{y}$} and the regressors $\{x_{it}\}$ into the rows of $\mathbf{X}$} yields ${\mathbf{y}}={\mathbf{X}}\boldsymbol{\beta}+{\mathbf{Z}}\boldsymbol{\eta}+\boldsymbol{\epsilon}$, where $\boldsymbol{\eta}$ collects the individual effects $c_{i}$ and $\mathbf{Z}$} is the $n\times N$ incidence matrix (with $N$ the number of individuals) that has a one in the row corresponding to observation $(i,t)$ and column $i$ when that observation belongs to individual $i$. In this formulation the assignment structure encoded by $\mathbf{Z}$} may itself be stochastic and potentially correlated with the random effects $\boldsymbol{\eta}$, which is precisely the type of dependence considered by Karl & Zimmerman (2021). In the conventional RE panel-data model, the exogeneity condition $E[c_{i}\mid x_{i1},\ldots,x_{iT_{i}}]=0$ restricts the dependence between the random effects $c_{i}$ and the regressors $x_{it}$. By contrast, the zero-bias conditions listed in the previous subsection include a requirement that rules out dependence between the random effects and the random-effects design matrix $\mathbf{Z}$}. In simple random-intercept panels $\mathbf{Z}$} is just an incidence matrix and is often implicitly assumed to be fixed without further consideration.
Rather than testing a global null that the RE estimator is consistent, the bias diagnostic of Karl & Zimmerman (2021) quantifies the finite-sample bias in individual estimable functions $\mathbf{k}^{\prime}\boldsymbol{\beta}$ that arises when the random-effects design matrix $\mathbf{Z}$} and the random-effects vector $\boldsymbol{\eta}$ are dependent, with $\mathbf{X}$} treated as fixed. The plug-in quantity $\hat{\boldsymbol{\nu_{k}}}^{\prime}\hat{\boldsymbol{\eta}}$ and its permutation-based $p$-value can therefore be viewed as a parameter-specific analogue of a Hausman-type specification check, obtained from a single fitted mixed model. In applications where both $\mathbf{X}$} and $\mathbf{Z}$} may reasonably be regarded as stochastic, the diagnostic still targets the component of bias that operates through the joint behavior of $\mathbf{Z}$} and $\boldsymbol{\eta}$; it does not attempt to separate any additional bias that might arise from stochastic variation in $\mathbf{X}$} alone. In this sense, the Karl & Zimmerman (2021) diagnostic focuses on a narrower class of misspecification than the classical Hausman or Mundlak–Wooldridge tests, and is best viewed as a complementary, parameter-specific tool rather than a replacement for those global specification checks.
An advantage of this approach is that it can be implemented directly from a single fitted mixed model, without requiring the analyst to construct a full FE or CRE analogue in settings that depart from the standard individual-by-time panel, such as models with multiple-membership random effects, nontrivial error-covariance structures $\mathbf{R}$, or a very large number of random-effect levels. In such applications, FE or CRE-style specifications are often still possible in principle, but may require additional data construction and modeling work. The value-added model in Section 4 provides one such example. When conventional FE or CRE estimators are readily available, we recommend using the classical Hausman or Mundlak–Wooldridge tests as global checks of the RE specification, and then using the bias diagnostic to identify which coefficients or contrasts exhibit the largest finite-sample bias and in which direction.
3. Gasoline Consumption Data Application
We illustrate both tests using the Gasoline dataset from the R package plm, which contains panel data on gasoline consumption (Baltagi & Griffin, 1983; Croissant & Millo, 2008). Listing 1 presents the R code to run the Hausman test and to load libraries for the bias test. All results in this section can be reproduced using Listings 1 and 2 together with the Gasoline data supplied by the plm package.
1# Load required libraries
2library(plm)
3library(lme4)
4library(mixedbiastest)
5# Load the Gasoline dataset
6data("Gasoline", package = "plm")
7# Define the model formula
8form <- lgaspcar ~ lincomep + lrpmg + lcarpcap
9# Within (FE) model
10wi <- plm(form, data = Gasoline, model = "within")
11# Random effects model
12re <- plm(form, data = Gasoline, model = "random")
13# Hausman test
14htest <- phtest(wi, re)
15# Results
16print(htest)
Listing 1:  Hausman Test for the Gasoline Dataset
Hausman Test
chisq = 302.8, df = 3, p-value < 2.2e-16
The significant result rejects the null hypothesis that the RE estimator is consistent.
Next, we fit the mixed model with lme4 and run the bias diagnostic (Listing 2). For comparison of parameter estimates, we also fit the corresponding fixed effects model, although this is not required by the bias test. Although the random-effects estimator from plm and the random-intercept estimator from lmer are not exactly identical, they are very similar in this simple random-intercept setting. We therefore use plm for the Hausman test and lmer for the bias diagnostic, since mixedbiastest is implemented on top of lme4.
1# mixed model
2random_model <- lmer(lgaspcar ~ lincomep + lrpmg + lcarpcap + (1 | country),
3 data = Gasoline)
4# run the bias test
5res<-mixedbiastest(random_model, n_permutations = 1e6, verbose = FALSE)
6# treat country indicators as fixed effects
7fixed_model <- lm(lgaspcar ~ lincomep + lrpmg + lcarpcap + country,
8 data = Gasoline)
9print(random_model)
10print(fixed_model)
11print(res)
Listing 2:  Bias Test for the Gasoline Dataset
[…]
Table 1.  Gasoline Data: FE, RE Estimates, RE - FE difference, Internal Bias Estimates, and Bias p-value (rounded)
Parameter FE Est RE Est (RE - FE) Bias Estimate Bias p-value
(Intercept) 2.29 2.15 -0.14 -0.17 0.1048
lincomep 0.66 0.59 -0.07 -0.04 0.1596
lrpmg -0.32 -0.37 -0.05 -0.04 0.0008
lcarpcap -0.64 -0.62 0.02 0.01 0.1975
The graphical summaries of the bias estimates and reference distributions appear in Figure 1. The results from Listing 2 are combined into Table 1. The Hausman test rejects the null hypothesis that the RE estimator is consistent, and the bias diagnostic suggests that the RE estimator for lrpmg is downward biased. Note the similarity of the internal bias estimates from the RE model to the RE–FE differences in Table 1.
4. Value-added Model (VAM) Results with GPvam
Linear mixed models are used for value-added assessment of teachers based on the standardized test scores of their students (Morganstein & Wasserstein, 2014). There are many variations of these models, but one particular instance is the complete persistence (CP) VAM (Ballou et al., 2004; Mariano et al., 2010; Karl et al., 2013b). In the CP model, intra-student correlation is modeled with off-diagonal entries in the error covariance matrix ($\mathbf{R}$). Classroom heterogeneity – the vector of individual “teacher effects” (Lockwood et al., 2007) – is modeled with random effects ($\boldsymbol{\eta}$). A different variance component is fit for teachers in different grades, producing a diagonal covariance matrix $\mathbf{G}$} with the number of unique elements equal to the number of grades in the study. The random-effects model matrix $\mathbf{Z}$} tracks the assignment of students to classrooms, and also attributes student scores after the first year to their previous years’ teachers, producing a “multiple-membership” structure for $\mathbf{Z}$} (Browne et al., 2001). The nonrandom assignment of students to classrooms creates the potential for bias (Lockwood & McCaffrey, 2007) according to the mechanism described by Karl & Zimmerman (2021), through dependence between the random-effects design matrix $\mathbf{Z}$} and the teacher effects $\boldsymbol{\eta}$.
These features make the CP VAM a natural setting in which to use a parameter-specific bias diagnostic. The random-effects structure involves thousands of teacher effects, the student-level error covariance matrix $\mathbf{R}$ is block-diagonal rather than homoskedastic, and the multiple-membership design encoded by $\mathbf{Z}$} is more complex than a simple random-intercept panel. The Karl & Zimmerman (2021) diagnostic can be applied directly to the fitted mixed model and returns bias estimates and permutation-based $p$-values for individual fixed effects and user-specified contrasts, without requiring any change to the original random-effects specification.
The R package GPvam (Karl et al., 2012, 2013b) in Listing 3 fits the CP model and has a function to calculate the bias diagnostic. Note that GPvam requires the provided R data object to contain columns named “teacher”, “year”, and “student” so that the $\mathbf{Z}$}, $\mathbf{G}$}, and $\mathbf{R}$ matrices can be constructed. This is implicit in the function call, and the formula argument to GPvam only constructs the fixed-effects design matrix $\mathbf{X}$} and parameter vector $\boldsymbol{\beta}$. We examine the bias results for the application of the CP model to math test scores from a large urban elementary school dataset with 2834 students from grades 4 to 6. This is the same data set used by Karl et al. (2013a). In this value-added model application, we use the GPvam implementation of the bias diagnostic, which accommodates the block-diagonal error covariance matrix and multiple-membership random effects structure that fall outside the scope of the current mixedbiastest implementation.
1library(GPvam)
2result.cp<-GPvam(data, persistence = "CP",
3 formula(~as.factor(Race_Ethnicity)+as.factor(year)+as.factor(Gender)+0),
4 hessian = TRUE,tol1=1e-9)
5# Perform the bias test on all fixed effects
6bias.test.custom(result.cp)
7# Perform the bias test on a contrast: White - Hispanic
8bias.test.custom(result.cp,n_perms=1e6, k_vectors = list(c(0,0,0,-1,1,0,0,0)))
Listing 3:  Fitting a CP model with GPvam and performing the bias diagnostic
Table 2.  CP Model Fixed Effects Estimates, Standard Errors, Internal Bias Estimates, and Bias p-values
Fixed Effect Estimate Std. Error Bias Estimate Bias p-value
Black 24.2894 0.1612 0.0062 0.7430
Am Indian/Alaska 24.3683 0.3228 -0.0364 0.1570
Asian/Pac Island 25.7910 0.1863 0.0506 0.0105
Hispanic 24.3989 0.1058 -0.0691 0.0004
White 25.4836 0.0894 0.0597 0.0000
year=1 - - - -
year=2 0.9831 0.0600 0.0023 0.5440
year=3 1.9738 0.0876 0.0012 0.7530
Gender=F - - - -
Gender=M 0.0552 0.0819 0.0036 0.4990
The CP estimates and bias diagnostics for the fixed effects appear in Table 2. The diagnostic indicates that the estimator for the Hispanic effect is downward biased, while the estimators for the White and Asian/Pac Island effects are upward biased. The diagnostic for the contrast White - Hispanic reveals even stronger bias: the estimated contrast is $0.1287$, and none of the one million permutations exceeded the test statistic ($p$-value $\approx 0$). Although the bias is statistically significant, its magnitude may not be practically significant for this particular application. Nevertheless, this example illustrates how the diagnostic can facilitate the investigation of potential bias due to nonrandom allocation of random effects on a case-by-case basis. In this example we do not propose a change in estimator based solely on the diagnostic. Instead, the bias estimates are used descriptively to flag coefficients and contrasts whose interpretation may be sensitive to nonrandom teacher assignment.
5. Conclusion
We have illustrated how the bias test of Karl & Zimmerman (2021) can be used to detect instances where the mixed model (1) produces biased estimators for fixed effects. Besides considering bias rather than consistency, this test is narrower than the Hausman test in the sense that it focuses on bias due to dependence between the random effects and their design matrix $\mathbf{Z}$} in situations where the fixed- and random-effects design matrices are not orthogonal (with respect to the inverse of the error covariance matrix). In contrast, the Hausman test is sensitive to any potential source of inconsistency in the random-effects estimator of the fixed effects, provided the fixed-effects estimator is consistent. Nevertheless, the bias diagnostic can be computed from the fitted random-effects model without changing the original RE specification and generates summaries for each fixed effect or contrast. These features make it a potential complement to the Hausman test when assessing potential misspecification in linear mixed models.
In applications where both RE and FE (or CRE) estimators are readily available, we view the Hausman or Mundlak–Wooldridge tests as primary specification checks: a rejection of the null typically leads the analyst to prefer an estimator that is robust to correlation between regressors and random effects. The bias diagnostic is then informative about which coefficients or contrasts are most affected and by how much. When the conclusions from the two approaches differ, we recommend using the diagnostic in a sensitivity-analysis role rather than as a replacement for the classical specification tests. A convenient workflow is therefore to (i) fit the RE model, (ii) apply a FE- or CRE-based Hausman/Mundlak–Wooldridge test, and (iii) when the test is rejected or borderline, use the mixed-model bias diagnostic to identify and report the coefficients or contrasts that are most affected.