### v5

# Parameter-Specific Bias Diagnostics in Random-Effects Panel Data Models

Abstract.
The Hausman specification test assesses the random-effects specification by comparing the random-effects estimator with a fixed-effects alternative. This note shows how a recently proposed bias diagnostic for linear mixed models can complement that test in random-effects panel-data applications. The diagnostic delivers parameter-specific internal estimates of finite-sample bias, together with permutation-based $p$-values, from a single fitted random-effects model. We illustrate its use in a gasoline-demand panel and in a value-added model for teacher evaluation using publicly available R packages, and we discuss how the resulting coefficient-specific bias summaries can be incorporated into routine practice.
Key words and phrases:
linear mixed model, Hausman test, correlated random effects, permutation test
1. Introduction
Consider a linear mixed model:
(1) ${\mathbf{y}}={\mathbf{X}}\boldsymbol{\beta}+{\mathbf{Z}}\boldsymbol{\eta}+\boldsymbol{\epsilon},$
with $\boldsymbol{\eta}\sim N(\boldsymbol{0},{\mathbf{G}})$ independent of $\boldsymbol{\epsilon}\sim N(\boldsymbol{0},\mathbf{R})$. Here $\mathbf{y}$} is an $n\times 1$ response vector, $\mathbf{X}$} is an $n\times p$ fixed-effects design matrix with associated coefficient vector $\boldsymbol{\beta}$, and $\mathbf{Z}$} is an $n\times m$ random-effects design matrix with random-effects vector $\boldsymbol{\eta}$. The matrices $\mathbf{G}$} and $\mathbf{R}$ denote the $m\times m$ random-effects covariance matrix and the $n\times n$ error covariance matrix, respectively.
In random-effects panel-data applications, consistency of the RE estimator requires the unobserved individual effects to be uncorrelated (or, more generally, mean independent) with the regressors; see, for example, Hausman & Taylor, 1981a. The Hausman test (Hausman, 1978) is the classical tool for assessing this specification by comparing the random-effects (RE) and fixed-effects (FE) estimators. In its original formulation, this comparison involves re-fitting the model with $\boldsymbol{\eta}$ treated as an additional vector of fixed effects, although correlated random-effects (CRE) implementations in the spirit of Mundlak (1978) and Wooldridge (2010) deliver Hausman-type specification tests from a single augmented RE fit; see Section 2. Under the null hypothesis, both RE and FE are consistent, so their estimates should differ only by sampling variation. Under the alternative, only the FE estimator remains consistent. A large discrepancy is therefore evidence against the RE specification (Hausman, 1978; Hausman & Taylor, 1981b).
The Hausman test concerns consistency: it asks whether the RE estimator converges to the true parameter values as the sample size grows. Bias, by contrast, is the systematic difference between an estimator’s expectation and the target parameter at a given sample size. These are distinct concepts – an estimator may be asymptotically consistent yet still exhibit finite-sample bias – but similar forms of misspecification, especially failures of exogeneity, can affect both. It is therefore natural to use a bias diagnostic alongside the Hausman test when assessing possible misspecification. In what follows, we keep these roles separate: the Hausman test serves as a classical large-sample specification check for the RE estimator, whereas the diagnostic of Karl & Zimmerman (2021) is used as a finite-sample tool to summarize the magnitude and direction of parameter-specific bias within the same RE fit.
We draw on the work of Karl & Zimmerman (2021), who derive an internal diagnostic for bias in linear mixed models when, conditional on the fixed-effects design matrix $\mathbf{X}$}, the random-effects design matrix $\mathbf{Z}$} is stochastic and may be dependent on the random effects $\boldsymbol{\eta}$. Their diagnostic produces a plug-in bias estimate, $\hat{\boldsymbol{\nu_{k}}}^{\prime}\hat{\boldsymbol{\eta}}$, and a permutation-based $p$-value without requiring a second, fixed-effects-only model fit. This feature is especially useful in complex applications (such as the value-added model in Section 4), where treating $\boldsymbol{\eta}$ as fixed would be impractical because of insufficient degrees of freedom or the absence of off-the-shelf software. For the theoretical development of the bias formula and the permutation scheme, see Karl & Zimmerman (2021).
In this note, we show how the bias diagnostic can be used alongside the Hausman specification test. The Hausman test is global and asymptotic, whereas the internal bias diagnostic is parameter-specific and finite-sample. The quantity $\hat{\boldsymbol{\nu_{k}}}^{\prime}\hat{\boldsymbol{\eta}}$ is large when the fitted random effects align strongly with the model-dependent weighting vector $\hat{\boldsymbol{\nu_{k}}}$; in the stochastic-$\mathbf{Z}$} setting of Karl & Zimmerman, 2021, that alignment is induced by dependence between $\mathbf{Z}$} and $\boldsymbol{\eta}$. Our aim is not to propose a replacement for the Hausman or Mundlak–Wooldridge tests, but rather to show how an existing mixed-model bias diagnostic can supplement them by providing coefficient- or contrast-specific information from a single fitted mixed model.
Section 2 reviews the Hausman test and the bias diagnostic. Section 3 applies these ideas to a gasoline-consumption panel using the plm and mixedbiastest packages, and Section 4 illustrates the diagnostic in a value-added modeling (VAM) application using GPvam.
2. Hausman and Bias Diagnostics
2.1. The Hausman Test
The Hausman test (Hausman, 1978) compares the FE and RE estimators for $\boldsymbol{\beta}$, $\hat{\boldsymbol{\beta}}_{\text{FE}}$ and $\hat{\boldsymbol{\beta}}_{\text{RE}}$. Under the null hypothesis, both estimators are consistent. Letting $\mathbf{M}$$^{-}$ denote a generalized inverse of a matrix $\mathbf{M}$}, the test statistic
$H=(\hat{\boldsymbol{\beta}}_{\text{RE}}-\hat{\boldsymbol{\beta}}_{\text{FE}})^{\prime}\bigl[\operatorname{Var}(\hat{\boldsymbol{\beta}}_{\text{FE}})-\operatorname{Var}(\hat{\boldsymbol{\beta}}_{\text{RE}})\bigr]^{-}(\hat{\boldsymbol{\beta}}_{\text{RE}}-\hat{\boldsymbol{\beta}}_{\text{FE}})$
has an asymptotic $\chi^{2}(r)$ null distribution, where $r$ is the rank of $\operatorname{Var}(\hat{\boldsymbol{\beta}}_{\text{FE}})-\operatorname{Var}(\hat{\boldsymbol{\beta}}_{\text{RE}})$ (Hausman & Taylor, 1981a). A significant value of $H$ provides evidence against the RE orthogonality assumptions and, under the usual conditions, against consistency of the RE estimator. In the canonical panel-data setting, this statistic is derived under homoskedastic, serially uncorrelated idiosyncratic errors, that is, $\mathbf{R}=\sigma^{2}\mathbf{I}_{n}$. In the linear mixed model formulation we write the model with a general error covariance matrix $\mathbf{R}$; the gasoline example below maintains the standard homoskedastic assumption, whereas the value-added model in Section 4 uses a block-diagonal $\mathbf{R}$.
2.2. Bias Diagnostic
Let $\hat{{\mathbf{G}}}$ and $\hat{\mathbf{R}}$ denote estimates of $\mathbf{G}$} and $\mathbf{R}$, and define $\hat{{\mathbf{V}}}={\mathbf{Z}}\hat{{\mathbf{G}}}{\mathbf{Z}}^{\prime}+\hat{\mathbf{R}}$. The empirical random-effects (RE) estimator of $\boldsymbol{\beta}$ is
(2) $\hat{\boldsymbol{\beta}}_{\text{RE}}=({\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{X}})^{-}{\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{y}}.$
To connect directly with Karl & Zimmerman (2021), we first present the diagnostic under their setup, which conditions on $\mathbf{X}$} and treats $\mathbf{Z}$} as stochastic. For any $p\times 1$ vector $\mathbf{k}$ such that $\mathbf{k}^{\prime}\boldsymbol{\beta}$ is an estimable linear combination of the fixed effects (for example, a single coefficient $\beta_{j}$ or a contrast such as $\beta_{j}-\beta_{\ell}$), define the $1\times m$ row vector
(3) $\hat{\boldsymbol{\nu_{k}}}^{\prime}\equiv\mathbf{k}^{\prime}({\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{X}})^{-}{\mathbf{X}}^{\prime}\hat{{\mathbf{V}}}^{-1}{\mathbf{Z}}.$
Under the conditions of Theorem 2 of Karl & Zimmerman (2021),
(4) $\operatorname{Bias}(\mathbf{k}^{\prime}\hat{\boldsymbol{\beta}}_{\text{RE}})=\mathbf{E}[\hat{\boldsymbol{\nu_{k}}}^{\prime}\boldsymbol{\eta}].$
Thus the finite-sample bias of the RE estimator for $\mathbf{k}^{\prime}\boldsymbol{\beta}$ is governed by the alignment between the random effects $\boldsymbol{\eta}$ and the coefficient-specific weighting vector $\hat{\boldsymbol{\nu_{k}}}$. Plugging the empirical best linear unbiased predictor $\hat{\boldsymbol{\eta}}$ into the right-hand side of (4) yields the internal bias estimate $\hat{\boldsymbol{\nu_{k}}}^{\prime}\hat{\boldsymbol{\eta}}$. This bias is zero, for example, if (i) ${\mathbf{X}}^{\prime}\mathbf{R}^{-1}{\mathbf{Z}}=\boldsymbol{0}$ with probability one, which implies $\hat{\boldsymbol{\nu_{k}}}=\boldsymbol{0}$, or (ii) in the stochastic-$\mathbf{Z}$} framework of Karl & Zimmerman (2021), $\boldsymbol{\eta}$ is independent of $\mathbf{Z}$}, which implies $\mathbf{E}[\hat{\boldsymbol{\nu_{k}}}^{\prime}\boldsymbol{\eta}]=0$. A third way to remove this source of bias is to treat $\boldsymbol{\eta}$ as fixed (that is, fit an FE model), in which case the corresponding FE estimator of $\mathbf{k}^{\prime}\boldsymbol{\beta}$ remains unbiased under the same $\boldsymbol{\epsilon}\perp{\mathbf{Z}}$ condition, provided that $\mathbf{k}^{\prime}\boldsymbol{\beta}$ remains estimable (Karl & Zimmerman, 2021).
[…]
2.3. Relation to correlated random effects and Mundlak–Wooldridge tests
In the panel-data literature, a common way to test the random-effects specification is through correlated random effects (CRE) or “Mundlak” type augmentations of the RE model. Suppose we have the panel-data setup
$y_{it}=x_{it}^{\prime}\beta+c_{i}+\varepsilon_{it},$
with individual effect $c_{i}$ and idiosyncratic error $\varepsilon_{it}$. Under the usual random-effects assumptions, including strict exogeneity and $E[c_{i}\mid x_{i1},\ldots,x_{iT_{i}}]=0$ (often summarized as $\operatorname{Cov}(x_{it},c_{i})=0$), the random-effects estimator is consistent. Mundlak (1978) models potential correlation between $x_{it}$ and $c_{i}$ by including the group means $\bar{x}_{i}$ in the regression. Wooldridge (2010) shows that a simple Hausman-type alternative is to estimate the augmented random-effects model
$y_{it}=x_{it}^{\prime}\beta+\bar{x}_{i}^{\prime}\gamma+c_{i}+\varepsilon_{it},$
and apply a (robust) Wald test of $H_{0}:\gamma=0$. This CRE test uses only an augmented RE fit but is asymptotically equivalent to the traditional FE–RE Hausman test when the idiosyncratic errors are homoskedastic and serially uncorrelated (Hausman & Taylor, 1981a). Robust implementations based on covariance estimators that are consistent under general forms of heteroskedasticity and serial correlation are widely used in applied work; see, e.g., Wooldridge (2010).
In terms of the linear mixed model (1), stacking the panel outcomes $\{y_{it}\}$ into $\mathbf{y}$} and the regressors $\{x_{it}\}$ into the rows of $\mathbf{X}$} yields ${\mathbf{y}}={\mathbf{X}}\boldsymbol{\beta}+{\mathbf{Z}}\boldsymbol{\eta}+\boldsymbol{\epsilon}$, where $\boldsymbol{\eta}$ collects the individual effects $c_{i}$ and $\mathbf{Z}$} is the $n\times N$ incidence matrix (with $N$ the number of individuals) that has a one in the row corresponding to observation $(i,t)$ and column $i$ when that observation belongs to individual $i$. In a conventional random-intercept panel, this incidence matrix is usually treated as fixed once the sample has been defined. The classical RE exogeneity condition then concerns dependence between the random effects $c_{i}$ and the regressors $x_{it}$, whereas the formal stochastic-$\mathbf{Z}$} framework of Karl & Zimmerman (2021) concerns dependence between $\boldsymbol{\eta}$ and the assignment structure encoded by $\mathbf{Z}$}. The two setups are therefore related but not identical.
For standard panels, we view $\hat{\boldsymbol{\nu_{k}}}^{\prime}\hat{\boldsymbol{\eta}}$ primarily as a descriptive, coefficient-specific summary from the fitted RE model rather than as a literal test of a stochastic-$\mathbf{Z}$} assignment mechanism. Because $\hat{\boldsymbol{\nu_{k}}}$ depends on the fitted RE weighting for the coefficient or contrast of interest, large values of $\hat{\boldsymbol{\nu_{k}}}^{\prime}\hat{\boldsymbol{\eta}}$ often accompany sizeable FE–RE differences for the same coefficient. In applications where the assignment structure encoded by $\mathbf{Z}$} is itself plausibly stochastic – for example, multiple-membership or assignment problems – the interpretation developed by Karl & Zimmerman (2021) applies more directly. In either case, the diagnostic is best viewed as complementary to, rather than a substitute for, the Hausman or Mundlak–Wooldridge tests: it does not test the global null $E[c_{i}\mid x_{i1},\ldots,x_{iT_{i}}]=0$, but instead supplies a parameter-specific finite-sample summary from a single fitted mixed model.
An advantage of this approach is that it can be implemented directly from a single fitted mixed model, without requiring the analyst to construct a full FE or CRE analogue in settings that depart from the standard individual-by-time panel, such as models with multiple-membership random effects, nontrivial error-covariance structures $\mathbf{R}$, or a very large number of random-effect levels. In such applications, FE or CRE-style specifications are often still possible in principle, but may require additional data construction and modeling work. The value-added model in Section 4 provides one such example. When conventional FE or CRE estimators are readily available, we recommend using the classical Hausman or Mundlak–Wooldridge tests as global checks of the RE specification, and then using the bias diagnostic to identify which coefficients or contrasts exhibit the largest finite-sample bias and in which direction.
3. Gasoline Consumption Data Application
We illustrate both diagnostics using the Gasoline dataset from the plm package, which contains panel data on gasoline consumption (Baltagi & Griffin, 1983; Croissant & Millo, 2008). Listing 1 shows the code for the Hausman test, and Listing 2 shows the code for the bias diagnostic. All results in this section can be reproduced using those two listings together with the Gasoline data supplied by plm.
1# Load required libraries
2library(plm)
3library(lme4)
4library(mixedbiastest)
5# Load the Gasoline dataset
6data("Gasoline", package = "plm")
7# Define the model formula
8form <- lgaspcar ~ lincomep + lrpmg + lcarpcap
9# Within (FE) model
10wi <- plm(form, data = Gasoline, model = "within")
11# Random effects model
12re <- plm(form, data = Gasoline, model = "random")
13# Hausman test
14htest <- phtest(wi, re)
15# Results
16print(htest)
Listing 1:  Hausman Test for the Gasoline Dataset
Hausman Test
chisq = 302.8, df = 3, p-value < 2.2e-16
This overwhelmingly small $p$-value provides strong evidence against the RE specification for the slope coefficients.
Next, we fit the mixed model with lme4 and run the bias diagnostic (Listing 2). For comparison of coefficient estimates, we also fit the corresponding least-squares model with country dummies, although that second fit is not required by the diagnostic itself. The random-effects estimator from plm and the random-intercept estimator from lmer are not exactly identical, but they are very close in this simple random-intercept setting. We therefore use plm for the Hausman test and lmer for the bias diagnostic, since mixedbiastest is implemented on top of lme4.
1# mixed model
2random_model <- lmer(lgaspcar ~ lincomep + lrpmg + lcarpcap + (1 | country),
3 data = Gasoline)
4# run the bias diagnostic
5res<-mixedbiastest(random_model, n_permutations = 1e6, verbose = FALSE)
6# treat country indicators as fixed effects
7fixed_model <- lm(lgaspcar ~ lincomep + lrpmg + lcarpcap + country,
8 data = Gasoline)
9print(random_model)
10print(fixed_model)
11print(res)
Listing 2:  Bias Diagnostic for the Gasoline Dataset
[…]
Table 1.  Gasoline Data: FE and RE Estimates, RE - FE Difference, Internal Bias Estimates, and Permutation $p$-values (rounded)
Parameter FE Est RE Est (RE - FE) Bias Estimate Perm. $p$-value
(Intercept) 2.29 2.15 -0.14 -0.17 0.1048
lincomep 0.66 0.59 -0.07 -0.04 0.1596
lrpmg -0.32 -0.37 -0.05 -0.04 0.0008
lcarpcap -0.64 -0.62 0.02 0.01 0.1975
Remark. The “FE Est” intercept in Table 1 comes from the least-squares fit with country dummies and therefore depends on the dummy-variable normalization (choice of baseline category). It is not directly comparable to the intercept parameter in the random-intercept model, so we focus on the slope coefficients when interpreting FE–RE differences.
The graphical summaries of the internal bias estimates and reference distributions appear in Figure 1, and the numerical results from Listing 2 are summarized in Table 1. Because the country-incidence matrix is fixed once the sample is defined, we interpret these bias summaries descriptively rather than as evidence for a literal stochastic-$\mathbf{Z}$} assignment mechanism. On that descriptive reading, the diagnostic points most clearly to lrpmg: its internal bias estimate is negative and its permutation $p$-value is very small. Note also the similarity between the internal bias estimates from the RE model and the corresponding FE–RE differences in Table 1.
4. Value-added Model (VAM) Results with GPvam
Linear mixed models are widely used in value-added assessment of teachers based on their students’ standardized test scores (Morganstein & Wasserstein, 2014). One important instance is the complete persistence (CP) VAM (Ballou et al., 2004; Mariano et al., 2010; Karl et al., 2013b). In the CP model, intra-student correlation is represented through off-diagonal entries in the error covariance matrix $\mathbf{R}$. Between-classroom heterogeneity – the vector of individual “teacher effects” (Lockwood et al., 2007) – is modeled through random effects $\boldsymbol{\eta}$. A different variance component is fit for teachers in different grades, producing a diagonal covariance matrix $\mathbf{G}$} with one distinct diagonal element for each grade in the study. The random-effects model matrix $\mathbf{Z}$} records students’ classroom assignments and also attributes post-baseline scores to prior teachers, producing a multiple-membership structure for $\mathbf{Z}$} (Browne et al., 2001). Nonrandom assignment of students to classrooms can therefore create the sort of dependence between $\mathbf{Z}$} and $\boldsymbol{\eta}$ highlighted by Karl & Zimmerman (2021), and hence the potential for bias (Lockwood & McCaffrey, 2007).
These features make the CP VAM a natural setting for a parameter-specific bias diagnostic. The random-effects structure involves thousands of teacher effects, the student-level error covariance matrix $\mathbf{R}$ is block-diagonal rather than homoskedastic, and the multiple-membership design encoded by $\mathbf{Z}$} is more complex than a simple random-intercept panel. The Karl & Zimmerman (2021) diagnostic can be applied directly to the fitted mixed model and returns internal bias estimates and permutation-based $p$-values for individual fixed effects and user-specified contrasts, without requiring any change to the original random-effects specification.
The GPvam package (Karl et al., 2012, 2013b), illustrated in Listing 3, fits the CP model and includes a function for the bias diagnostic. The supplied R data object must contain columns named “teacher”, “year”, and “student”, because those variables determine the $\mathbf{Z}$}, $\mathbf{G}$}, and $\mathbf{R}$ matrices. This construction is implicit in the function call: the formula argument to GPvam specifies only the fixed-effects design matrix $\mathbf{X}$} and coefficient vector $\boldsymbol{\beta}$. We illustrate the method with mathematics test scores from a large urban elementary-school dataset containing 2834 students in grades 4 through 6, the same dataset analyzed by Karl et al. (2013a). In this application we use the GPvam implementation of the bias diagnostic because it accommodates the block-diagonal error covariance matrix and multiple-membership random-effects structure that lie outside the scope of the current mixedbiastest implementation.
1library(GPvam)
2result.cp<-GPvam(data, persistence = "CP",
3 formula(~as.factor(Race_Ethnicity)+as.factor(year)+as.factor(Gender)+0),
4 hessian = TRUE,tol1=1e-9)
5# Perform the bias diagnostic on all fixed effects
6bias.test.custom(result.cp)
7# Perform the bias diagnostic on a contrast: White - Hispanic
8bias.test.custom(result.cp,n_perms=1e6, k_vectors = list(c(0,0,0,-1,1,0,0,0)))
Listing 3:  Fitting a CP model with GPvam and performing the bias diagnostic
Table 2.  CP Model Fixed Effects Estimates, Standard Errors, Internal Bias Estimates, and Permutation $p$-values
Fixed Effect Estimate Std. Error Bias Estimate Perm. $p$-value
Black 24.2894 0.1612 0.0062 0.7430
Am Indian/Alaska 24.3683 0.3228 -0.0364 0.1570
Asian/Pac Island 25.7910 0.1863 0.0506 0.0105
Hispanic 24.3989 0.1058 -0.0691 0.0004
White 25.4836 0.0894 0.0597 0.0000
year=1 - - - -
year=2 0.9831 0.0600 0.0023 0.5440
year=3 1.9738 0.0876 0.0012 0.7530
Gender=F - - - -
Gender=M 0.0552 0.0819 0.0036 0.4990
The CP estimates and fixed-effect bias diagnostics appear in Table 2. The diagnostic suggests downward bias for the Hispanic coefficient and upward bias for the White and Asian/Pac Island coefficients. For the contrast White - Hispanic, the internal bias estimate is $0.1287$, and none of the one million sampled permutations produced a value as extreme, so the Monte Carlo $p$-value is effectively 0 at that simulation resolution. Although the statistical evidence for bias is strong, its practical importance depends on the scale on which the contrast will be interpreted. Even so, the example shows how the diagnostic can be used to investigate possible bias arising from nonrandom student assignment on a coefficient-by-coefficient basis. We do not propose changing estimators on the basis of the diagnostic alone; instead, we use the bias estimates descriptively to flag coefficients and contrasts whose interpretation may be sensitive to nonrandom teacher assignment.
5. Conclusion
We have illustrated how the diagnostic of Karl & Zimmerman (2021) can be used to flag fixed effects and contrasts for which a fitted mixed model may exhibit non-negligible finite-sample bias. Relative to the Hausman test, the diagnostic is narrower: it targets the coefficient-specific alignment term $\mathbf{E}[\hat{\boldsymbol{\nu_{k}}}^{\prime}\boldsymbol{\eta}]$ rather than every possible source of inconsistency in the RE estimator. Its formal sampling interpretation is sharpest in the stochastic-$\mathbf{Z}$} settings emphasized by Karl & Zimmerman (2021), where nonrandom assignment can induce alignment between the random-effects design and the random effects themselves. In more conventional panel settings, we view the resulting statistic primarily as a descriptive complement to FE–RE comparisons. Even so, the diagnostic is attractive because it can be computed from a single fitted RE model and reported for any fixed effect or contrast of substantive interest.
In applications where both RE and FE (or CRE) estimators are readily available, we view the Hausman or Mundlak–Wooldridge tests as primary specification checks: a rejection of the null typically leads the analyst to prefer an estimator that is robust to correlation between regressors and random effects. The bias diagnostic is then informative about which coefficients or contrasts are most affected and by how much. When the conclusions from the two approaches differ, we recommend using the diagnostic in a sensitivity-analysis role rather than as a replacement for the classical specification tests. A convenient workflow is therefore to (i) fit the RE model, (ii) apply a FE- or CRE-based Hausman/Mundlak–Wooldridge test, and (iii) when the test is rejected or borderline, use the mixed-model bias diagnostic to identify and report the coefficients or contrasts that are most affected.